# Generative Retrieval with Large Language Models

### v2

@@ 3.2.1 Page-level Results
Coarse-grained page-level results, as shown in Table 1, demonstrate that the LLM2GR framework, when implemented with Llama-2-13b, achieves the best R-precision scores of 57.77, 48.70, 83.69, and 57.63 on the NQ, HotpotQA, FEVER, and WoW datasets, respectively. This significantly surpasses the performance of sparse retrieval BM25 and dense retrieval Contriever in a zero-shot scenario. It also shows strong competitive power against the fully trained DPR method, especially on the WoW and FEVER datasets, with improvements of 27.08 and 31.01 points, respectively. This result is consistent with the hypothesis that LLMs are powerful in recalling coarse-grained title identifiers, enabling the acquisition of high-quality relevant pages that assist in the subsequent fine-grained recall stage.
@@ 3.2.2 Passage-level Results
Fine-grained reference passage results, as shown in Table 2, reveal that the LLM2GR framework, when implemented with Llama-2-13b, also achieves the best scores of 68.20, 30.04, 58.42, and 63.43 on the TriviaQA, HotpotQA, FEVER, and WoW datasets, respectively. We note that the improvement of the framework in fine-grained reference passage compared to the DPR method is relatively reduced compared to the page-level results. This suggests potential for optimization in activating LLMs to recall more detailed and longer reference, presenting a greater challenge compared to recalling shorter title. Notably, DPR performs excellently on the NQ dataset, which may relate to its training data format. Interestingly, in the HotpotQA dataset, BM25 remains competitive, surpassing dense retrieval methods, possibly due to the longer questions in this dataset leading to more vocabulary overlap. LLM2GR shows significant progress on the FEVER and WoW datasets, demonstrating the potential and adaptability of LLMs in recalling high-quality reference passage across different task formats. Furthermore, the general enhancement in performance with the progression from Llama to Llama-2 and the increase in model size indicates a correlation between the recall ability and the underlying capabilities of LLMs.
@@ 3.2.3 Downstream Task Results
Downstream task results are presented in Table 3. LLM2GR, based on Llama-2-13b recalled passage, achieved the best scores of 72.94, 78.79, and 14.77 on the TriviaQA, FEVER, and WoW downstream tasks, respectively, validating the performance of LLM recall references in downstream tasks. On the open-domain question answering NQ dataset, although DPR performed excellently after full data training, LLM2GR also displayed highly competitive performance. On the other hand, in the TriviaQA and HotpotQA datasets, due to the length of the questions, BM25 achieved excellent performance by obtaining more vocabulary overlap, yet LLM2GR still achieved comparable or better performance in most cases. The unsupervised trained Contriever performed relatively poorly across all tasks, emphasizing the crucial role of supervised training in enhancing the performance of dense retrieval models.
@@ 3.3 Ablation Study
Table 4: Ablation study results on the NQ, TriviaQA, and HotpotQA datasets are shown, with the left half showing the R-Precision at the coarse-grained page level and the right half showing Answer in Context for fine-grained passage. We compared the performance differences without weighted scores, without SPRL, and without the first stage.
In this subsection, we conduct ablation studies, comparing methods without weighted scores (w/o weight), without Short Prefix Recall Location (w/o SPRL), and without the first stage of document title recall (w/o first stage). The results are shown in Table 4.
Without weighted scores, relying solely on the recall scores from the second stage leads to a simultaneous decrease in performance for both coarse and fine-grained results, emphasizing the importance of considering scores from both stages. The model, by taking into account title scores, is more capable of selecting the correct document, and within the correct document, it is more likely to choose the correct reference. More results on the choice of weighted $\alpha$ can be found in Appendix D.
Without SPRL, recalling longer segments has a minor impact on page-level performance. However, it significantly affects the quality of fine-grained reference passage, where longer recall lengths paradoxically lead to decreased performance. This result is somewhat counterintuitive and might be due to all document knowledge being stored in the parameters during the pre-training phase, with a short prefix sufficient to locate the required reference. Longer references introduce redundancy and noise, thus lowering effectiveness. Notably, when using Llama-2-13b for recall, recalling complete passages on the NQ dataset takes about 600 minutes, while recalling short prefixes only requires 150 minutes, significantly reducing time costs. However, considering that dense retrieval takes about 20 minutes, further optimization of speed remains crucial. More experiments on prefix length can be found in Appendix E.
@@
3.4 Further Analysis
After General Fine-tuning of LLMs We also test the Vicuna model (Chiang et al., 2023) and the Llama-2-chat model refined through reinforcement learning from human feedback (Touvron et al., 2023b), both of which underwent general fine-tuning. This general fine-tuning did not significantly enhance the performance of LLMs in recalling and locating reference. This may be due to the paradigm difference between the fine-tuning data and the recall location task, coupled with the fact that most knowledge was already acquired during the pre-training phase. By creating more diverse recall instruction tuning data, further improvements in model performance might be achieved. Detailed results can be found in Appendix F.
Impact of Few-Shot We explore adding few-shot prompts in the second stage of fine-grained recall and observed its impact on overall performance. This approach brought slight improvements only in the HotpotQA dataset, while showing a slight decline in NQ and TriviaQA. Importantly, adding more few-shot examples significantly reduced generation speed. This suggests that, although few-shot prompting offers a potential path for improvement, extensive exploration is still needed to devise more effective prompting methods. Detailed results can be found in Appendix G.
Memory Usage Analysis Dense retrieval methods such as Contriever and DPR require over 60GB of memory usage. In contrast, sparse retrieval methods use far less memory, only needing 17GB. The LLM2GR framework, utilizing FM-index and Trie indexing, requires only 8GB when pre-encoding and storing all documents with FM-index, and the Trie storing all title identifiers needs just 25MB, which is negligible. Compared to sparse and dense retrieval methods, the recall framework effectively saves memory.
@@ 4 Related Work
Traditional methods of obtaining reference include sparse and dense retrieval. Sparse retrieval, using TF-IDF and BM25, matches questions and passages (Robertson et al., 2009; Chen et al., 2017; Yang et al., 2019). Recent approaches, such as ORQA (Lee et al., 2019) and DPR (Karpukhin et al., 2020), employ dense context vectors for passage indexing to enhance performance. However, in dual-encoder dense retrieval models, the representations of questions and passages are obtained independently, leading to performance limitations due to shallow vector interactions (Khattab and Zaharia, 2020).
Interest has surged in using autoregressive language models to generate identifiers to simplify the retrieval process and address the bottleneck of limited interactions in dual-encoder models. For example, Cao et al. (2021) generates Wikipedia page titles for retrieval, Tay et al. (2022) targets generating paths from root to leaf in hierarchical clustering trees, and Bevilacqua et al. (2022) maps to unique n-grams. Recently, Lee et al. (2022) proposed a generative multi-hop retrieval method, Li et al. (2023b) employed multiple identifiers to collaboratively determine retrieval passages, and Ren et al. (2023) introduced a two-stage approach that first generates passages, then URL identifiers. Ziems et al. (2023) utilized LLMs to first generate page URLs, followed by traditional retrieval techniques to obtain fine-grained passages. However, whether traditional or generative retrieval, they can only obtain predefined and segmented text chunks, making it difficult to naturally retrieve references from arbitrary positions.
Recent research has found that relevant knowledge can be extracted from LLMs through prompting, especially in domains with insufficient coverage in knowledge bases (Liu et al., 2022; Fang et al., 2022). Enhancing model performance through the output of LLMs has also gained attention. Liu et al. (2022); Sun et al. (2023); Yu et al. (2023) propose using GPT-3 to generate relevant context as references, treating these contexts as additional inputs when answering questions. However, fully generating context through LLMs is still plagued by the phenomenon of hallucination (Li et al., 2023a).
5 Conclusion and Future Work
This paper introduces a framework named LLM2GR, which utilizes LLMs to independently recall reference passage and can be flexibly applied to various knowledge-sensitive tasks. Mimicking the human habit of searching for information, we first prompt the LLM to recall relevant document pages, and then from these pages, recall and locate reference passage. Furthermore, through beam search constrained by Trie and FM-index, we ensure that the content recalled by the LLM is a subset of existing texts. This framework can be flexibly used with various open-source LLMs, expanding the potential applications of LLMs. In the future, we consider enhancing the ability of LLMs to recall under constraints through instruction tuning, as well as applying this framework to more retrieval fields, exploring lightweight methods to inject new document knowledge into LLMs, and integrating multi-hop reasoning into reference recall.
@@ 5 Conclusion and Future Work
Although LLM2GR demonstrates the potential of LLMs to recall reference passage in knowledge-sensitive tasks like humans, its application still faces several limitations. Firstly, this framework struggles to surpass the performance of the current state-of-the-art retrieval models, especially those models that have been fine-tuned on specific tasks through supervision. In the future, there is a need to explore more effective ways of instruction tuning for recalling under constraints. At the same time, LLM2GR relies on document title identifiers for phased recall, meaning that its recall capability may be limited for documents lacking clear titles or identifiers.
Moreover, the framework finds it challenging to effectively recall documents that appear less frequently in the pre-training stage. This indicates that if a document appears infrequently in the training data of the LLM, or if the document content significantly differs from the training data, LLM2GR may encounter difficulties in recalling these documents. For the updating of documents and the injection of new knowledge, LLM2GR requires additional training to inject this new information into the model parameters. There is still a need to explore more efficient, lightweight methods for injecting new documents in the future.
@@
Appendix A Constrained Decoding Methods
Figure 3: (a) Shows the process of a LLM generating title identifiers using a prefix tree. (b) Shows the process of a LLM generating passage prefixes in a document set via FM-index.
@@ Appendix A Constrained Decoding Methods
NQ Open Domain QA Question Extractive 2837
HotpotQA Open Domain QA Question Short Abstractive 5600
TriviaQA Open Domain QA Question Extractive 5359
ELI5 Open Domain QA Question Long Abstractive 1507
FEVER Fact Checking Claim Classification 10444
WoW Dialogue Conversation Long Abstractive 3054
@@
A.1 Trie
The overall process during constrained decoding using a Trie is shown in Figure 3(a). Taking the generation of the title “Testamentary Capacity" as an example, the LLM first selects “Testament" from the set of token strings that start all titles. Subsequently, we can obtain the set of token strings {and, ary} following the string “Testament". After the LLM selects “ary", we get the prefix string “Testamentary", and finally continue to select new strings from the next set of token strings until the end-of-sequence token </s>is encountered, ceasing generation.
A.2 FM-Index
The FM-index (Ferragina and Manzini, 2000) is a data structure used for text retrieval that can store text efficiently with linear space complexity and support fast substring search operations. It is constructed based on the Burrows-Wheeler Transform (BWT) (Burrows, 1994). BWT is a method that converts a string into a form that is easy to compress. Given a string, BWT produces a transformed string through the following steps: generate all cyclic shifts of the string, sort all these shifts lexicographically, take the last character of each sorted shifted string to form a new string, which is the BWT result. For example, for the string “CABAC", the process of building the FM-index is as follows:
$\begin{array}[]{cccccc}\mathbf{F}&&&&&\mathbf{L}\\ \$^{6}&C&A&B&A&C^{5}\\ A^{2}&B&A&C&\$&C^{1}\\ A^{4}&C&\$&C&A&B^{3}\\ B^{3}&A&C&\$&C&A^{2}\\ C^{5}&\$&C&A&B&A^{4}\\ C^{1}&A&B&A&C&\$^{6}\end{array}$
The overall process during constrained decoding using FM-index is shown in Figure 3(b). Considering the generated prefix “The Greece GDP warrants are not technically bonds as investors do" for example, it first starts from the string “The" generated from all corpus, and gets its corresponding L column string set {christ, Greece, Johan}. After “Greece" is selected by the LLM, we can get the next set {U, G, part}, and continue the iteration until reaching the set maximum prefix length to stop generating.
Appendix B Dataset Details
  • Natural Questions (NQ) (Kwiatkowski et al., 2019) is constructed from real anonymized aggregated queries submitted to the Google search engine, with answers being snippets from manually annotated Wikipedia articles.
  • TriviaQA (Joshi et al., 2017) comprises a set of questions and answers initially crawled from knowledge question-answering and quiz websites.
  • HotpotQA (Yang et al., 2018) contains a set of question-answer pairs based on Wikipedia, requiring multiple-step reasoning over multiple Wikipedia pages to answer each question.
  • ELI5 (Fan et al., 2019) is a large-scale corpus for long-form question answering, consisting of questions and answers from the Reddit forum “Explain Like I’m Five" (ELI5), which require detailed and in-depth responses to open-ended questions.
  • FEVER (Thorne et al., 2018) is one of the largest datasets for fact-checking, used to determine whether a statement is supported or refuted based on textual sources.
  • Wizard of Wikipedia (WoW) (Dinan et al., 2018) is a task that requires intelligent agents in open-domain dialogues to demonstrate knowledge usage. The dialogues are directly grounded in knowledge retrieved from Wikipedia.
All experiments are tested using the public validation set as divided in the official KILT. Additional details of the datasets are presented in Table 5. All the data used in this paper come from the KILT benchmark(Petroni et al., 2021), and KILT is MIT licensed [^3] https://opensource.org/licenses/MIT.
Appendix C Prompts
In this subsection, we introduce the prompts used in the first stage for recalling coarse-grained title identifiers, in the second stage for recalling fine-grained reference passage, and in downstream tasks.
C.1 Prompts for the First Stage
@@
LLM2GR(Llama-7b) 54.46 57.03 44.56 36.87 58.48 25.55
LLM2GR(Vicuna-1.3-7b) 48.47 47.99 40.79 35.28 56.41 23.75
LLM2GR(Llama-13b) 54.42 55.53 46.30 37.72 60.96 26.34
LLM2GR(Vicuna-1.3-13b) 52.73 46.61 43.41 36.55 67.29 26.63
LLM2GR(Llama-2-7b) 56.33 56.43 46.20 38.07 62.88 27.55
LLM2GR(Llama-2-chat-7b) 3.31 1.12 0.98 4.09 3.97 3.43
LLM2GR(Vicuna-1.5-7b) 50.76 51.73 41.23 34.16 55.98 24.43
LLM2GR(Llama-2-13b) 57.77 54.41 48.70 40.82 68.20 30.04
LLM2GR(Llama-2-chat-13b) 1.94 1.60 1.55 6.38 7.93 4.71
LLM2GR(Vicuna-1.5-13b) 52.24 56.34 45.90 37.22 63.24 27.14
@@
Figure 6: On the NQ, TriviaQA, and HotpotQA datasets, the page-level and passage-level experimental results for Llama-2-13b under {0,1,3,5}-shot few-shot prompts.
C.2 Prompts for the Second Stage
[…]
C.3 Prompts for Reading Comprehension
[…]
@@ C.3 Prompts for Reading Comprehension
Appendix D Experimental Results for Different Values of Alpha
In Figure 4, we compare the experimental results of LLM2GR when implemented based on Llama-2-13b with different $\alpha$ values. When $\alpha=0.0$, it’s equivalent to not having a two-stage weighted method, relying only on the scores from the second stage’s fine-grained passage recall, resulting in the selection of suboptimal reference. With the increase of $\alpha$, the model sees improvements in both page-level and passage-level results, proving the importance of the first stage document scores for the final reference selection. However, when $\alpha$ reaches 0.95 and continues to increase, the final performance actually decreases to some extent, indicating the need to find a balance between the two for better results.
Appendix E Experimental Results for Different Prefix Lengths
In Figure 5, we conduct experiments with LLM2GR recalling different numbers of prefix tokens based on Llama-2-13b. We observe that longer prefix lengths do not bring additional performance improvements; on the contrary, they lead to a decrease in performance. Existing LLMs still perform better when generating shorter passages under constraints; longer passages introduce additional noise, resulting in decreased performance. However, overly short prefixes might also lack sufficient information, leading to an inability to accurately select the desired passage as a reference.
Appendix F Experimental Results of LLMs After General Fine-tuning
In Table 6, we compare the model performance in recalling references after supervised fine-tuning (Vicuna-1.3 and Vicuna-1.5 (Chiang et al., 2023)) and reinforcement learning from human feedback (Llama-2-chat). It is observed that the Vicuna models, after supervised fine-tuning, do not show further improvements; in fact, performance slightly declines. This indicates that the memorization of document knowledge primarily occurs during the pre-training phase, and further enhancement may require specific fine-tuning data and generation methods designed for recalling under constraints. In contrast, models fine-tuned with reinforcement learning from human feedback significantly drop in performance, unable to effectively recall reference under constraints. We note that models trained with reinforcement learning from human feedback often start their outputs with polite phrases like “Sure!”, which affects the model distribution and thus leads to failure.
@@ Appendix F Experimental Results of LLMs After General Fine-tuning
Appendix G Experimental Results under Few-Shot Prompts
In Figure 6, we compare the experimental results of LLM2GR when incorporating few-shot prompts in the second stage of fine-grained passage recall based on Llama-2-13b. Using more sample prompts only brings partial improvements on the HotpotQA dataset, and the degree of improvement tends to stabilize when the number of samples increases from 3 to 5. On the NQ and TriviaQA datasets, no further improvements are observed, and a slight decline is even noted. There is still a need to explore more effective prompting methods.
Appendix H Experimental Results of Selecting Different Numbers of First Stage Documents

### v3

@@ 3.2.1 Page-level Results
Coarse-grained page-level results, as shown in Table 1, demonstrate that the MindRef framework, when implemented with Llama-2-13b, achieves the best R-precision scores of 57.77, 48.70, 83.69, and 57.63 on the NQ, HotpotQA, FEVER, and WoW datasets, respectively. This significantly surpasses the performance of sparse retrieval BM25 and dense retrieval Contriever in a zero-shot scenario. It also shows strong competitive power against the fully trained DPR method, especially on the WoW and FEVER datasets, with improvements of 27.08 and 31.01 points, respectively. This result is consistent with the hypothesis that LLMs are powerful in recalling coarse-grained title identifiers, enabling the acquisition of high-quality relevant pages that assist in the subsequent fine-grained recall stage.
@@ 3.2.2 Passage-level Results
Fine-grained reference passage results, as shown in Table 2, reveal that the MindRef framework, when implemented with Llama-2-13b, also achieves the best scores of 68.20, 30.04, 58.42, and 63.43 on the TriviaQA, HotpotQA, FEVER, and WoW datasets, respectively. We note that the improvement of the framework in fine-grained reference passage compared to the DPR method is relatively reduced compared to the page-level results. This suggests potential for optimization in activating LLMs to recall more detailed and longer reference, presenting a greater challenge compared to recalling shorter title. Notably, DPR performs excellently on the NQ dataset, which may relate to its training data format. Interestingly, in the HotpotQA dataset, BM25 remains competitive, surpassing dense retrieval methods, possibly due to the longer questions in this dataset leading to more vocabulary overlap. MindRef shows significant progress on the FEVER and WoW datasets, demonstrating the potential and adaptability of LLMs in recalling high-quality reference passage across different task formats. Furthermore, the general enhancement in performance with the progression from Llama to Llama-2 and the increase in model size indicates a correlation between the recall ability and the underlying capabilities of LLMs.
@@ 3.2.3 Downstream Task Results
Downstream task results are presented in Table 3. MindRef, based on Llama-2-13b recalled passage, achieved the best scores of 72.94, 78.79, and 14.77 on the TriviaQA, FEVER, and WoW downstream tasks, respectively, validating the performance of LLM recall references in downstream tasks. On the open-domain question answering NQ dataset, although DPR performed excellently after full data training, MindRef also displayed highly competitive performance. On the other hand, in the TriviaQA and HotpotQA datasets, due to the length of the questions, BM25 achieved excellent performance by obtaining more vocabulary overlap, yet MindRef still achieved comparable or better performance in most cases. The unsupervised trained Contriever performed relatively poorly across all tasks, emphasizing the crucial role of supervised training in enhancing the performance of dense retrieval models.
@@ 3.3 Ablation Study
Table 4: Ablation study results, with the left half showing the R-Precision at the coarse-grained page level and the right half showing Answer in Context for fine-grained passage.
@@
In this subsection, we compare methods without weighted scores (w/o weight), without Short Prefix Recall Location (w/o SPRL), and without the first stage of document title recall (w/o first stage). The results are shown in Table 4.
Without weighted scores, relying solely on the recall scores from the second stage leads to a simultaneous decrease in performance for both coarse and fine-grained results, emphasizing the importance of considering scores from both stages. The model, by taking into account title scores, is more capable of selecting the correct document, and within the correct document, it is more likely to choose the correct reference. More results on the choice of weighted $\alpha$ can be found in Figure 4.
Without SPRL, recalling longer segments has a minor impact on page-level performance. However, it significantly affects the quality of fine-grained reference passage, where longer recall lengths paradoxically lead to decreased performance. This result is somewhat counterintuitive and might be due to all document knowledge being stored in the parameters during the pre-training phase, with a short prefix sufficient to locate the required reference. Longer references introduce redundancy and noise, thus lowering effectiveness. Notably, when using LLaMA-2-13b for recall, recalling complete passages on the NQ dataset takes about 600 minutes, while recalling short prefixes only requires 150 minutes, significantly reducing time costs. However, considering that dense retrieval takes about 20 minutes, further optimization of speed remains crucial. More experiments on prefix length can be found in Figure 5.
@@ 4 Related Work
Traditional retrieval methods rely on sparse (TF-IDF, BM25 Robertson and Zaragoza (2009)) or dense (ORQA Lee et al. (2019), DPR Karpukhin et al. (2020)) representations. However, dual-encoder dense retrieval faces limitations due to shallow interactions between independently encoded question and passage representations Khattab and Zaharia (2010).
Recent work explores using LLMs to generate identifiers for retrieval, aiming to simplify the process and enhance interaction compared to dual-encoder models. These approaches target page titles De Cao et al. (2021), hierarchical paths Tay et al. (2022), n-grams Bevilacqua et al. (2022), multi-hop paths Lee et al. (2022), multiple identifiers Li et al. (2023b), or a two-stage approach with passages and URLs Ren et al. (2023); Yue et al. (2025). These methods, however, predominantly retrieve predefined text segments, hindering flexible retrieval from arbitrary positions within full documents.
Leveraging LLMs to directly generate relevant knowledge Fang et al. (2022) or augmenting models with LLM-generated context (e.g., GenRead Yu et al. (2023), A+B Tang et al. (2024)) has also shown promise for knowledge-intensive tasks. However, hallucination remains a significant challenge Li et al. (2023a), potentially providing unreliable or fabricated information.
5 Conclusion
This paper introduces MindRef, a framework utilizing LLMs to independently recall reference passages for knowledge-sensitive tasks. Mimicking human information-seeking behavior, the LLM first recalls relevant document pages, and then locates specific passages within them. Beam search, constrained by Trie and FM-index structures, ensures that recalled content is a subset of existing texts. This framework is adaptable to various open-source LLMs, broadening their potential applications.
@@ 5 Conclusion
Although MindRef demonstrates the potential of LLMs to recall reference passage in knowledge-sensitive tasks like humans, its application still faces several limitations. Firstly, this framework struggles to surpass the performance of the current SOTA retrieval models, especially those models that have been fine-tuned on specific tasks through supervision. In the future, there is a need to explore more effective ways of instruction tuning for recalling under constraints. At the same time, MindRef relies on document title identifiers for phased recall, meaning that its recall capability may be limited for documents lacking clear titles or identifiers.
Moreover, the framework finds it challenging to effectively recall documents that appear less frequently in the pre-training stage. This indicates that if a document appears infrequently in the training data of the LLM, or if the document content significantly differs from the training data, MindRef may encounter difficulties in recalling these documents. For the updating of documents and the injection of new knowledge, MindRef requires additional training to inject this new information into the model parameters. There is still a need to explore more efficient, lightweight methods for injecting new documents in the future.
@@
Appendix A Preliminary
A.1 Trie
(a) Title Generation with Prefix Tree.
(b) Passage Prefix Generation with FM-index.
Figure 3: Constrained Decoding Methods: (a) Shows the process of an LLM generating title identifiers using a prefix tree. (b) Shows the process of an LLM generating passage prefixes in a document set via FM-index.
The overall process during constrained decoding using a Trie is shown in Figure 3(a). Taking the generation of the title "Testamentary Capacity" as an example, the LLM first selects "Testament" from the set of token strings that start all titles. Subsequently, we can obtain the set of token strings {and, ary} following the string "Testament". After the LLM selects "ary", we get the prefix string "Testamentary", and finally continue to select new strings from the next set of token strings until the end-of-sequence token </s>is encountered, ceasing generation.
A.2 FM-Index
The FM-index (Ferragina and Manzini, 2000) is a data structure used for text retrieval that can store text efficiently with linear space complexity and support fast substring search operations. It is constructed based on the Burrows-Wheeler Transform (BWT) (Burrows et al., 1994). BWT is a method that converts a string into a form that is easy to compress. Given a string, BWT produces a transformed string through the following steps: generate all cyclic shifts of the string, sort all these shifts lexicographically, take the last character of each sorted shifted string to form a new string, which is the BWT result. For example, for the string "CABAC", the process of building the FM-index is as follows:
$\begin{array}[]{cccccc}\mathbf{F}&&&&&\mathbf{L}\\ \mathdollar^{6}&C&A&B&A&C^{5}\\ A^{2}&B&A&C&\mathdollar&C^{1}\\ A^{4}&C&\mathdollar&C&A&B^{3}\\ B^{3}&A&C&\mathdollar&C&A^{2}\\ C^{5}&\mathdollar&C&A&B&A^{4}\\ C^{1}&A&B&A&C&\mathdollar^{6}\end{array}$
The overall process during constrained decoding using FM-index is shown in Figure 3(b). Considering the generated prefix "The Greece GDP warrants are not technically bonds as investors do" for example, it first starts from the string "The" generated from all corpus, and gets its corresponding L column string set {christ, Greece, Johan}. After "Greece" is selected by the LLM, we can get the next set {U, G, part}, and continue the iteration until reaching the set maximum prefix length to stop generating.
Appendix B Additional Details for Experiments
B.1 Datasets
@@
NQ (Kwiatkowski et al., 2019) Open-domain QA Question Extractive 2837
HotpotQA (Yang et al., 2018) Open-domain QA Question Short Abstractive 5600
TriviaQA (Joshi et al., 2017) Open-domain QA Question Extractive 5359
ELI5 (Fan et al., 2019) Open-domain QA Question Long Abstractive 1507
FEVER (Thorne et al., 2018) Fact Checking Claim Classification 10444
WoW (Dinan et al., 2019) Dialogue Conversation Long Abstractive 3054
@@
We conduct extensive experiments on 6 knowledge-sensitive tasks from the KILT benchmark (Petroni et al., 2021). These tasks include open-domain QA tasks such as NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), and ELI5 (Fan et al., 2019), the fact-checking task FEVER (Thorne et al., 2018), and the open-domain dialogue system WoW (Dinan et al., 2019). All experiments are tested using the public validation set as divided in the official KILT. Additional details of the datasets are presented in Table 5. All the data used in this paper come from the KILT benchmark(Petroni et al., 2021), and KILT is MIT licensed [^4] https://opensource.org/licenses/MIT. We evaluate the quality of coarse-grained pages and fine-grained reference passage, as well as the enhancement of reference for downstream tasks.
B.2 Prompt
In this subsection, we introduce the prompt used in the first stage for recalling coarse-grained title identifiers, in the second stage for recalling fine-grained reference passage, and in downstream tasks.
B.2.1 Prompt for the First Stage
[…]
B.2.2 Prompt for the Second Stage
[…]
B.2.3 Prompt for Reading Comprehension
[…]
@@ B.2.3 Prompt for Reading Comprehension
MindRef (LLaMA-7b) 54.46 57.03 44.56 36.87 58.48 25.55
MindRef (Vicuna-1.3-7b) 48.47 47.99 40.79 35.28 56.41 23.75
MindRef (LLaMA-13b) 54.42 55.53 46.30 37.72 60.96 26.34
MindRef (Vicuna-1.3-13b) 52.73 46.61 43.41 36.55 67.29 26.63
MindRef (LLaMA-2-7b) 56.33 56.43 46.20 38.07 62.88 27.55
MindRef (LLaMA-2-chat-7b) 3.31 1.12 0.98 4.09 3.97 3.43
MindRef (Vicuna-1.5-7b) 50.76 51.73 41.23 34.16 55.98 24.43
MindRef (LLaMA-2-13b) 57.77 54.41 48.70 40.82 68.20 30.04
MindRef (LLaMA-2-chat-13b) 1.94 1.60 1.55 6.38 7.93 4.71
MindRef (Vicuna-1.5-13b) 52.24 56.34 45.90 37.22 63.24 27.14
@@
Figure 6: On the NQ, TriviaQA, and HotpotQA datasets, the page-level and passage-level experimental results for LLaMA-2-13b under {0,1,3,5}-shot few-shot prompt.
@@
B.3 Further Analysis
Different Values of Alpha In Figure 4, we compare the experimental results of MindRef when implemented based on Llama-2-13b with different $\alpha$ values. When $\alpha=0.0$, it’s equivalent to not having a two-stage weighted method, relying only on the scores from the second stage’s fine-grained passage recall, resulting in the selection of suboptimal reference. With the increase of $\alpha$, the model sees improvements in both page-level and passage-level results, proving the importance of the first stage document scores for the final reference selection. However, when $\alpha$ reaches 0.95 and continues to increase, the final performance actually decreases to some extent, indicating the need to find a balance between the two for better results.
Different Prefix Lengths In Figure 5, we conduct experiments with MindRef recalling different numbers of prefix tokens based on Llama-2-13b. We observe that longer prefix lengths do not bring additional performance improvements; on the contrary, they lead to a decrease in performance. Existing LLMs still perform better when generating shorter passages under constraints; longer passages introduce additional noise, resulting in decreased performance. However, overly short prefixes might also lack sufficient information, leading to an inability to accurately select the desired passage as a reference.
After General Fine-tuning of LLMs. We also test the Vicuna model (Chiang et al., 2023) and the LLaMA-2-chat model refined through reinforcement learning from human feedback (Touvron et al., 2023b), both of which underwent general fine-tuning. This general fine-tuning did not significantly enhance the performance of LLMs in recalling and locating reference. This may be due to the paradigm difference between the fine-tuning data and the recall location task, coupled with the fact that most knowledge was already acquired during the pre-training phase. By creating more diverse recall instruction tuning data, further improvements in model performance might be achieved. Detailed results can be found in Table 6.
Impact of Few-Shot. We explore adding few-shot prompt in the second stage of fine-grained recall and observed its impact on overall performance. This approach brought slight improvements only in the HotpotQA dataset, while showing a slight decline in NQ and TriviaQA. Importantly, adding more few-shot examples significantly reduced generation speed. This suggests that, although few-shot prompting offers a potential path for improvement, extensive exploration is still needed to devise more effective prompting methods. Detailed results can be found in Figure 6.